# Random pairing MLE for estimation of item parameters in Rasch model

### v1

@@ 3 Main results
In this section, we collect the main theoretical guarantees for $\mathsf{RP\text{-}MLE}$ and its variant $\mathsf{MRP\text{-}MLE}$. Section 3.1 focuses on the finite sample $\ell_{\infty}$ error bound. While one can translate $\ell_{\infty}$ error into an $\ell_{2}$ bound, in Section 3.2, we present a much sharper characterization of the $\ell_{2}$ error of $\mathsf{RP\text{-}MLE}$. Lastly in Section 3.3, we provide a distributional characterization of $\mathsf{RP\text{-}MLE}$.
3.1 $\ell_{\infty}$ estimation error and top-$K$ recovery
Without loss of generality, we assume that the scores of the items are ordered, i.e., $\theta_{1}^{\star}\geq\theta_{2}^{\star}\geq\cdots\geq\theta_{m}^{\star}$, and denote $\Delta_{K}\coloneqq\theta_{K}^{\star}-\theta_{K+1}^{\star}$. In words, $\Delta_{K}$ measures the difference between the difficulty levels of items $K$ and $K+1$. The following theorem provides $\ell_{\infty}$ error bounds and top-$K$ recovery guarantee for both $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$. We defer its proof to Section 4.2.
@@ 3.1 $\ell_{\infty}$ estimation error and top-$K$ recovery
Suppose that $mp\geq 2$ and $np\geq C_{1}\kappa_{1}^{4}\kappa_{2}^{5}\log^{3}(n)$ for some sufficiently large constant $C_{1}>0$. Suppose that there exists some constant $\alpha>0$ such that $m\leq n^{\alpha}$. Let $\widehat{\bm{\theta}}$ be the $\mathsf{RP\text{-}MLE}$ estimator. With probability at least $1-O(n^{-10})$, $\widehat{\bm{\theta}}$ satisfies
@@
Here $C_{2},C_{3}>0$ are some universal constants. All the claims continue to hold for $\mathsf{MRP\text{-}MLE}$ as long as there exists some constant $\beta>0$ such that $n_{\mathrm{split}}\leq n^{\beta}$.
@@
Intuitively, as we have $\Theta(mnp)$ sampled comparisons and $m$ parameters to estimate, the optimal $\ell_{\infty}$ error should be $O(\sqrt{m/(mnp)})=O(1/\sqrt{np})$. This is formalized in the following result from [NZ23].
Proposition 1 (Minimax lower bound, Theorems 3.3 and 3.4 in [NZ23]).
Assume that $np\geq C_{1}$ for some suffiently large constant $C_{1}>0$. For any $n$ and $m$, there exists a class of user and item parameters $\Theta$ such that
$\inf_{\hat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{E}\left\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\right\|_{2}^{2}\geq\frac{C_{2}m}{np},\qquad\text{and}\qquad\inf_{\hat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{E}\left\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\right\|_{\infty}^{2}\geq\frac{C_{2}}{np},$
@@
$\inf_{\hat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{P}\left[\widehat{\bm{\theta}}\text{ fails to identify all top-}K\text{ items}\right]\geq\frac{1}{2}.$
Comparing our upper bounds with the lower bound in the proposition, we can see that both $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ are rate-optimal in $\ell_{\infty}$ estimation error and top-$K$ recovery sample complexity, up to logarithmic and $\kappa$ factors.
@@
Second, it is a standard and necessary requirement to have $np\gtrsim\log(n)$ to make sure that each item is compared to at least one user with high probability. In Theorem 1 we require an extra $\log^{2}(n)$ factor to suppress a quadratic error term that comes up in the analysis. This cubic log factor can possibly be loose, but it is a minor issue and we leave it as a problem for future research.
@@
Comparison with [NZ23].
The closest result to our paper in terms of $\ell_{\infty}$ guarantee for the Rasch model appears in the recent work [NZ23]. Their spectral method uses a similar construction of the item-item comparisons but without disjoint pairing. To provide detailed comparisons, we restate their results below.
Proposition 2 (Informal, Theorem 3.1 in [NZ23]).
@@
This error rate is similar to ours. However, the required sample size is much larger as they require $p\gtrsim\log(m)/\sqrt{n}$. Our result makes a significant improvement by allowing a much smaller sampling rate $p$, cf. $mp\geq 2$ and $np\gtrsim\log^{3}(n)$. In fact, as we have argued earlier, it is nearly the sparsest possible. In addition, our methods enjoy a significantly better error rate dependency on $\kappa$. In Section 5.1, we provide empirical evidence for this improvement: when $\kappa$ is large, $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ outperform the spectral methods in [NZ23].
3.2 Refined $\ell_{2}$ error characterization
@@
The $\ell_{\infty}$ error guarantee in Theorem 1 immediately implies an $\ell_{2}$ error bound
@@
which is rate-optimal compared to the minimax lower bound in Proposition 1. However, this guarantee is only correct in an order-wise sense. In this section, we present a refined characterization of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ that is precise in the leading constant.
We start with necessary notation. Let $z_{ij}\coloneqq e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}$ and $\widehat{z}_{ij}\coloneqq e^{\hat{\theta}_{i}}e^{\hat{\theta}_{j}}/(e^{\hat{\theta}_{i}}+e^{\hat{\theta}_{j}})^{2}$. Define a matrix
$\bm{L}_{L\tilde{z}}=\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}\widetilde{z}_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$ (4)
for $\tilde{z}\in\{z,\widehat{z}\}$ and let $\bm{L}_{L\tilde{z}}^{\dagger}$ be its pseudo-inverse. One can view $\bm{L}_{L\tilde{z}}$ as a weighted graph Laplacian. In the following theorem, we show that $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ concentrates tightly around $[\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})]^{1/2}$. We defer the analysis to Section 4.3 and the complete proof to Section D.

### v2

@@ 3 Main results
In this section, we collect the main theoretical guarantees for $\mathrm{RP\text{-}MLE}$ and its variant $\mathrm{MRP\text{-}MLE}$. Section 3.1 focuses on the finite sample $\ell_{\infty}$ error bound. In Section 3.2, we present a non-asymptotic expansion that describes the distribution of $\mathrm{RP\text{-}MLE}$, and we apply this expansion to reach a Berry-Esseen type theorem and a much sharper characterization of the $\ell_{2}$ error of $\mathrm{RP\text{-}MLE}$. Lastly in Section 3.3, we prove the asymptotic normality of $\mathrm{MRP\text{-}MLE}$ when $m$ and $p$ are fixed and draw a connection between $\mathrm{MRP\text{-}MLE}$ and (weighted) pseudo MLE.
3.1 $\ell_{\infty}$ error bounds and top-$K$ recovery
Without loss of generality, we assume that the scores of the items are ordered, i.e., $\theta_{1}^{\star}\geq\theta_{2}^{\star}\geq\cdots\geq\theta_{m}^{\star}$, and denote $\Delta_{K}\coloneqq\theta_{K}^{\star}-\theta_{K+1}^{\star}$. In words, $\Delta_{K}$ measures the difference between the difficulty levels of items $K$ and $K+1$. The following theorem provides $\ell_{\infty}$ error bounds and top-$K$ recovery guarantee for both $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$. We defer its proof to Section A.2.
@@ 3.1 $\ell_{\infty}$ error bounds and top-$K$ recovery
Suppose that $mp\geq 2$ and $np\geq C_{1}\kappa_{1}^{4}\kappa_{2}^{5}\log^{3}(n)$ for some sufficiently large constant $C_{1}>0$. Suppose that there exists some constant $\alpha>0$ such that $m\leq n^{\alpha}$. Let $\widehat{\bm{\theta}}$ be the $\mathrm{RP\text{-}MLE}$ estimator. With probability at least $1-O(n^{-10})$, $\widehat{\bm{\theta}}$ satisfies
@@
Here $C_{2},C_{3}>0$ are some universal constants. All the claims continue to hold for $\mathrm{MRP\text{-}MLE}$ as long as there exists some constant $\beta>0$ such that $n_{\mathrm{split}}\leq n^{\beta}$.
@@
Based on results from the ranking literature, it is reasonable to guess that the optimal $\ell_{\infty}$ error is $O(\sqrt{m/(mnp)})=O(1/\sqrt{np})$. Here $m$ is the number of item parameters, and mnpmnp is the number of user-item comparisons. This guess is indeed correct, as formalized in the following result from NZ (23).
Proposition 1 (Minimax lower bound, Theorems 3.3 and 3.4 in NZ (23)).
Assume that $np\geq C_{1}$ for some sufficiently large constant $C_{1}>0$. For any $n$ and $m$, there exists a class of user and item parameters $\Theta$ such that
$\inf_{\widehat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{E}\left\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\right\|_{2}^{2}\geq\frac{C_{2}m}{np},\qquad\text{and}\qquad\inf_{\widehat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{E}\left\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\right\|_{\infty}^{2}\geq\frac{C_{2}}{np},$
@@
$\inf_{\widehat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{P}\left[\widehat{\bm{\theta}}\text{ fails to identify all top-}K\text{ items}\right]\geq\frac{1}{2}.$
Comparing our upper bounds with the lower bound in the proposition, we can see that both $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ are rate-optimal in $\ell_{\infty}$ estimation error and top-$K$ recovery sample complexity, up to logarithmic and $\kappa$ factors.
@@
Second, it is a standard and necessary requirement to have $np\gtrsim\log(n)$ to make sure that each item is compared to at least one user with high probability. In Theorem 1 we require an extra $\log^{2}(n)$ factor to suppress a quadratic error term that comes up in the analysis. This cubic log factor can possibly be loose, but it is a minor issue and we leave it to future research.
@@
Comparison with NZ (23).
The closest result to our paper in terms of $\ell_{\infty}$ guarantee for the Rasch model appears in the recent work by NZ (23). Their spectral method uses a similar construction of the item-item comparisons but without disjoint pairing. To provide detailed comparisons, we restate their results below.
Proposition 2 (Informal, Theorem 3.1 in NZ (23)).
@@
This error rate is similar to ours. However, the required sample size is much larger as they require $p\gtrsim\log(m)/\sqrt{n}$. Our result makes a significant improvement by allowing a much smaller sampling rate $p$, cf. $mp\geq 2$ and $np\gtrsim\log^{3}(n)$. In fact, as we have argued earlier, it is nearly the sparsest possible regime for estimating item parameters. In addition, our methods enjoy a significantly better error rate dependency on $\kappa$. In Section 4.1, we provide empirical evidence for this improvement: when $\kappa$ is large, $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ outperform the spectral methods in NZ (23).
Analysis via reduction to BTL model.
As mentioned in Section 2.2, the random pairing in $\mathrm{RP\text{-}MLE}$ reduces the problem to the Bradley-Terry-Luce model with non-uniform sampling. This reduction results in an item-item comparison graph with nice spectral properties, and allows us to invoke the general theory of MLE in the BTL model established in the recent work by YCOM (24). See Section A.2 for the full analysis.
3.2 Non-asymptotic expansion of $\mathrm{RP\text{-}MLE}$
An important aspect of statistical estimators is the quantification of the variability. In this section we provide a non-asymptotic expansion, which precisely characterizes of the distribution of the $\mathrm{RP\text{-}MLE}$ estimator $\widehat{\bm{\theta}}$. We supplement this result with a Berry-Esseen theorem and an application to obtain a precise $\ell_{2}$ error characterization.
We start with some necessary notation. Let $\sigma(x)=e^{x}/(1+e^{x})$ be the sigmoid function. Let $z_{ij}\coloneqq e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}$, $\widehat{z}_{ij}\coloneqq e^{\hat{\theta}_{i}}e^{\hat{\theta}_{j}}/(e^{\hat{\theta}_{i}}+e^{\hat{\theta}_{j}})^{2}$ and $\epsilon_{ij}^{t}\coloneqq Y_{ji}^{t}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})$. Let $L_{\mathrm{total}}\coloneqq\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}$ be the total number of observed comparisons in $\mathcal{G}_{Y}$.
We define $\bm{\bm{B}}\in\mathbb{R}^{m\times L_{\mathrm{total}}}$ and $\widehat{\bm{\epsilon}}\in\mathbb{R}^{L_{\mathrm{total}}}$ via
$\bm{B}\coloneqq\left[\cdots,\sqrt{z_{ij}}(\bm{e}_{i}-\bm{e}_{j}),\cdots\right]_{i>j:(i,j)\in\mathcal{E}_{Y}}\text{$\quad$(repeat $L_{ij}$ times for edge $(i,j)$)}$
and
$\widehat{\bm{\epsilon}}\coloneqq\left[\cdots,\epsilon_{ij}^{t}/\sqrt{z_{ij}},\cdots\right]_{(i,j,t):i>j,(i,j)\in\mathcal{E}_{Y},L_{ij}^{t}=1}\in\mathbb{R}^{L_{\mathrm{total}}}.$
Moreover, define a weighted graph Laplacian
$\bm{L}_{L\tilde{z}}\coloneqq\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}\widetilde{z}_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$ (3)
for $\tilde{z}$ being $z$ or $\widehat{z}$, and let $\bm{L}_{L\tilde{z}}^{\dagger}$ be its pseudo-inverse.
The following theorem characterizes the distribution of $\mathrm{RP\text{-}MLE}$ with a non-asymptotic expansion. The analysis is deferred to Section A.3 and the full proof is deferred to Section C.1.
Theorem 2.
Instate the assumptions of Theorem 1. With probability at least $1-O(n^{-10})$, the estimator $\widehat{\bm{\theta}}$ given by the Algorithm 1 can be written as
$\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ $=-\left[\nabla^{2}\mathcal{L}(\bm{\theta}^{\star})\right]^{\dagger}\nabla\mathcal{L}(\bm{\theta}^{\star})+\bm{r}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r},$ (4)
This theorem shows $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ can be well approximated by $\nabla^{2}\mathcal{L}(\bm{\theta}^{\star})]^{\dagger}\nabla\mathcal{L}(\bm{\theta}^{\star})$, a form that frequently appears in the analysis of maximum likelihood estimators. Moreover, it can be written as a linear transformation of the random vector $\widehat{\bm{\epsilon}}$, whose entries are independent outcomes of the item-item comparisons shifted and scaled to be mean-zero and variance-one. The term $\bm{L}_{Lz}^{\dagger}\bm{B}$ accounts for the geometry induced by the comparison graph $\mathcal{G}_{Y}$. As the residual term $\bm{r}$ has small magnitude, we may analyze the properties of $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ by focusing on the leading term $\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$.
We compare our result with the inference result in CLOX (23). Theorem 8 therein studies the Rasch model and provides inferential results for a joint estimator of $\bm{\theta}^{\star}$ and $\bm{\zeta}^{\star}$. However, it requires a dense sampling scheme when $n\geq m$, with
$p\gtrsim\sqrt{\frac{1}{m}}\vee\frac{\log^{2}(m)}{n}\vee\frac{n\log^{2}(n)}{m^{2}}.$
It also requires that both $m$ and $n$ tend to infinity. These two assumptions are significantly more restrictive than ours.
Normal approximation.
The main term in (4) can be approximated with a normal random variable, allowing various applications such as hypothesis testing on $\bm{\theta}^{\star}$. Formally, we present the following Berry-Esseen type theorem. The proof is deferred to Section C.2.
Proposition 3.
Instate the assumptions of Theorem 1. Let $\bm{x}$ be a normal random variable in $\mathbb{R}^{m}$ with variance $\bm{L}_{Lz}^{\dagger}$. Let $\mathcal{C}_{m}$ be the set of all the measurable convex subset of $\{\bm{\theta}\in\mathbb{R}^{m}:\bm{\theta}^{\top}\bm{1}_{m}=0\}$. Then we have that
$\sup_{A\in\mathcal{C}_{m}}\left|\mathbb{P}\left[\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\in A\mid\mathcal{G}_{Y}\right]-\mathbb{P}(\bm{x}\in A)\right|\leq C_{1}n^{-10}+C_{2}\frac{m^{5/4}\kappa_{1}^{3/2}\kappa_{2}^{3/2}}{(np)^{1/2}}.$
Refined $\ell_{2}$ error characterization.
@@ 3.2 Non-asymptotic expansion of $\mathrm{RP\text{-}MLE}$
Another possible application of Theorem 2 is a refined characterization of the $\ell_{2}$ estimation error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$. The $\ell_{\infty}$ error guarantee in Theorem 1 immediately implies an $\ell_{2}$ error bound
@@
which is rate-optimal compared to the minimax lower bound in Proposition 1. However, this guarantee is only correct in an order-wise sense. Here, we present a refined characterization of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ that is precise in the leading constant. In the following theorem, we show that $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ concentrates tightly around $\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})]^{1/2}$. We defer the complete proof to Section C.3.
Proposition 4.
$\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}\right|$ $\leq C_{1}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{\log(n)}{np}}+\frac{C_{2}\kappa_{1}^{6}\sqrt{m}\log^{2}(n)}{np};$ (6)
Theorem 4 is more refined compared to (5). First, it provides both upper and lower bounds for $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$. Second, there is no hidden constant in front of the leading term $\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})]^{1/2}$. In addition, inspecting the proof of Theorem 4, we see that
and the same holds for $\mathrm{Trace}(\bm{L}_{L\widehat{z}}^{\dagger})]^{1/2}$. Consequently, the right hand sides of both (6) and (7) are lower order terms compared to $\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})]^{1/2}$ when $n,m\rightarrow\infty$. Indeed this recovers the naive $\ell_{2}$ bound (5) under appropriate sample size assumptions.
Analysis via projected gradient descent trajectory.
Inspired by Che (23), we analyze MLE via the projected gradient descent trajectory. This approach gives us an alternative to the leave-one-out type argument in CFMW (19); CGZ (22) and the leave-two-out argument in GSZ (23). In particular, unlike the leave-one-out and leave-two-out arguments, the projected gradient descent approach does not require independence in the sampling of the compared pairs. This is crucial to our analysis as the disjoint pairing in Step 1(a) of Algorithm 1 induces dependent item-item edges. See Section A.3 for the full analysis.
3.3 Asymptotic normality of $\mathrm{MRP\text{-}MLE}$ and pseudo MLE
In this section, we consider the inference setting where $m$ and $p$ are fixed and $n$ tends to infinity. We establish the asymptotic normality of $\mathrm{MRP\text{-}MLE}$ and connect it to a weighted variant of pseudo MLE, which we call $\mathrm{WP\text{-}MLE}$. We show that these two estimators are asymptotically equal in distribution.
We start by describing the setup of an infinite sequence of estimators $\{\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}\}$ and relevant notations. We will consider the user parameters to be fixed and study the asymptotic normality of $\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}$ that accounts for the randomness in sampling, random pairing, and the comparison between paired outcomes. We use subscripts to denote the source of randomness in each expectation. For sampling, i.e., the generation of comparison graph $\mathcal{G}_{X}$, we use $\mathbb{E}_{\mathrm{s}}$; for random pairing, where we match item-item-user tuple, we use $\mathbb{E}_{\mathrm{r}}$; for forming an item-item comparison, i.e., determining whether the responses $X_{ti}$ and $X_{tj}$ are different for a given item-item-user tuple $(i,j,t)$, we use $\mathbb{E}_{\mathrm{d}}$; for comparison, i.e., given $X_{ti}\neq X_{tj}$, whether $X_{ti}<X_{tj}$, we use $\mathbb{E}_{\mathrm{c}}$. Multiple letters can be combined with + sign in the subscript. Let $\mathcal{L}_{k}(\bm{\theta})$ be the loss function for $\mathrm{RP\text{-}MLE}$ using $k$-th random splitting and $\mathcal{L}_{k}^{(t)}(\bm{\theta})$ be the sum of the terms corresponding to user $t$, i.e., $\mathcal{L}_{k}(\bm{\theta})\coloneqq\sum_{t=1}^{n}\mathcal{L}_{k}^{(t)}(\bm{\theta})$ and
$\mathcal{L}_{k}^{(t)}(\bm{\theta})\coloneqq-\sum_{\begin{subarray}{c}(i,j):i>j:\\ (i,j,t)\in\Omega_{k}\end{subarray}}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right].$
Here $\Omega_{k}$ denotes the set of paired item-item-user tuples for the $k$-th splitting.
We assume there is an infinite sequence of users, which has an infinite sequence of user parameters $\{\zeta_{n}^{\star}\}_{n\in\mathbb{N}^{+}}$, random sampling $\{A_{it}:i\in[m]\}_{t=1}^{\infty}$, and responses $\{X_{it}:A_{it}=1\}_{t=1}^{\infty}$. Moreover we assume there are random splittings $\Omega_{k}$ for $k=1,\ldots,n_{\mathrm{split}}$. We label the estimators as $\widehat{\bm{\theta}}_{(k)}^{(n)}$ to denote the $\mathrm{MRP\text{-}MLE}$ with $k$-th splitting and the random sampling, responses, and random splittings associated with the first $n$ users. Similarly we use $\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}$ to denote $\mathrm{WP\text{-}MLE}$ with first $n$ users. Moreover, we suppose that the following limits of average expectation and covariance matrices exist:
$\bm{H}^{\infty}$ $\coloneqq\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla^{2}\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star});$ (8a)
$\bm{V}_{\mathrm{same}}^{\infty}$ $\coloneqq\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})^{\top};$ (8b)
$\bm{V}_{\mathrm{diff}}^{\infty}$ $\coloneqq\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})^{\top}$ (8c)
Note that by the Cauchy-Schwarz inequality (see Section D.4 for a complete proof), we have
$\bm{V}_{\mathrm{diff}}^{\infty}\preceq\bm{V}_{\mathrm{same}}^{\infty}.$ (9)
The two sides are the same when $m_{t}=2$ for all $t$.
With this setup of a infinite sequence of estimators, we may study the asymptotic normality of $\mathrm{MRP\text{-}MLE}$ as $n$ tends to infinity, accounting for all randomness in $\mathbb{E}_{\mathrm{s+r+d+c}}$. We have the following result on $\mathrm{MRP\text{-}MLE}$ $\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}=(1/n_{\mathrm{split}})\sum_{i=1}^{n_{\mathrm{split}}}\widehat{\bm{\theta}}_{(i)}^{(n)}$. The analysis is deferred to Section A.4 and the full proof is deferred to Section D.1.